OpenStudy (anonymous):

Solve 15 ≥ -3x or x ≥ -2. {x | x ≤ -5} {x | x ≥ -5} {all reals}

OpenStudy (austinl):

\[15\ge-3x\] \[\frac{15}{-3}\le\frac{-3x}{-3}\] \[-5\le x\]

OpenStudy (anonymous):

thnank you so much

OpenStudy (austinl):

Do you know why the sign flipped?